OpenStudy (anonymous):

Find an equation of the largest sphere with center (6, 10 , 8) that is contained completely in the first octant.

OpenStudy (anonymous):

if centerd @(h,k,l)=(6,10,8) (x-6)^2+(y-10)^2+(z-8)^2=r^2

OpenStudy (anonymous):

if centered @ origin use x^2 +y^2 +z^2 = r^2

OpenStudy (anonymous):

radius cannot be larger than 6 else outside octant so (x-6)^2+(y-10)^2+(z-8)^2=6^2